The choices are listed in the screen shot below the questions. Only answer is needed, solving steps are not required. Thank you!

The returns of two portfolios were recorded for fourteen years. Portfolio 1 had a variance of returns of 155, while portfolio1.4085 0.6387 1.2513 1.5657 0.7992The p-value associated with this test statistic is0.6921 0.5702 0.2149 0.4119 0.4298Based on your results, what is your conclusion? O Reject the null hypothesis, and conclude that the variation in portfolio 1-1.5494 0.5256 0.5026 0.4704 0.60764.8771 4.9093 4.6634 4.6807 4.0346
Answer #1

(first part)

An F distribution with 13 and 13 degrees of freedom

(second part)

1.5657

F=s12/s22=155/99=1.5657

(third part)

0.2149

p-value=0.2149

(fourth part)

Do not reject the null hypothesis, and conclude that there is not enough evidence to claim that the variation in portfolio 1

Do not reject null hypothesis, and conclude that there is not enough .........

As the p-value=0.2149 is more than level of significance alpha=0.05(5%)

(fifth part)  LCL=0.5026 and UCL=4.877

4.8771 0.5026

100(1-alpha ) percent confidence interval for sigma1-squared /sigma2-squared is

        V

LCL=1.5657/3.115=0.5026

UCL==1.5657/0.311=4.877